Design and draw voltage regulator to provide 4 to 10 V variable output voltage using IC 7812.
1
Expert's answer
2020-07-03T10:05:01-0400


circuit diagram will look like as above.


Equation for finding output voltage is as

"R_2 = R_1 \frac{1}{\frac{V_{IC}}{V_{out}}-1}"

or it can be written as

"V_{out} = \frac{V_{IC}R_2}{R_1+R_2}"


Value of "V_{IC} = 12V"

Now for obtaining the output voltage 4-10 V we need to choose R1 and R2 such that we get the output.


If "R_2 = 5*R_1" then Vout = 10V


If "R_2 = 0.5R_1" then Vout = 4V